Find the area that is inside r=4-2cos∅ and the outside r=6+2cos∅.


1
Expert's answer
2021-04-25T17:14:26-0400

"r=4-2\cos \theta" and "r=6+2\cos \theta"

Let us find points of intersection:

"4-2\cos \theta =6+2\cos \theta"

"4\cos \theta=-2"

"\cos \theta =-\frac{1}{2}"

"\theta _1=\frac{2\pi}{3}" and "\theta _2=\frac{4\pi}{3}"


Formula of area of the region bounded by the graph of "r=f(\theta )" between the radial lines "\theta =\alpha" and "\theta =\beta" is "A=\frac{1}{2}\int\limits_{\alpha }^{\beta}f(\theta)^2d\theta" . Using this formula, we obtain

 "A=\frac{1}{2}\int\limits _{\frac{2\pi}{3}}^{\frac{4\pi}{3}}\left((4-2\cos \theta)^2-(6+2\cos \theta)^2
\right)d\theta=\frac{1}{2}\int\limits _{\frac{2\pi}{3}}^{\frac{4\pi}{3}}\left(-20-40\cos \theta
\right)d\theta ="

"=-10\int\limits _{\frac{2\pi}{3}}^{\frac{4\pi}{3}}\left(1+2\cos \theta
\right)d\theta =-10\left(\theta+2\sin \theta\right)\bigg|^\frac{4\pi}{3}_{\frac{2\pi}{3}}="

"=-10\left(\frac{4\pi}{3}-2\cdot \frac{\sqrt{3}}{2}\right)+10\left(\frac{2\pi}{3}+2\cdot \frac{\sqrt{3}}{2}\right)=20\sqrt{3}-\frac{20\pi}{3}"



Answer: "A= 20\sqrt{3}-\frac{20\pi}{3}" .